# Congestion-Aware Path Re-routing Strategy for Dense Urban Airspace

### v1

@@ IV-B Stream $X$ queuing system
        If the descend condition is not satisfied ($b_{k}=0$) with probability $\rho C(0)B(0)$, then the arrival $a_{k}$ will enter the $\beta_{Z}$ queue.
        ${}^{\beta}v_{1,k}=a_{k}\ ,$ ${}^{\beta}V_{1}(z)=A(z)$ (57)
        ${}^{\beta}v_{j+1,k+1}={}^{\beta}v_{j,k}+a_{k}\ ,$ ${}^{\beta}V_{j+1}(z)={}^{\beta}V_{j}(z)A(z)$ (58)
  2. 4.
    If the system is idle ($\theta_{k}=1$) and there are no UAS with higher preference ($a_{\mathrm{I,k}}=0$ , ${}^{\gamma}v_{L-1,k}=0$ ) with probability $\rho$ then
    $\rho=(1-\Theta_{0})A_{\mathrm{I}}(0){}^{\gamma}V_{L-1}(0)$ (59)
    1. (a)
      If there is no conflict because $a_{k}=0$, $c_{k}=\{0,1\}$ with probability $\rho A(0)$, then $c_{k}$ enters service. No UAS enters the $\beta_{Z}$ queue.
      ${}^{\beta}v_{1,k}=0\ ,$ ${}^{\beta}V_{1}(z)=1$ (60)
      ${}^{\beta}v_{j+1,k+1}={}^{\beta}v_{j,k}\ ,$ ${}^{\beta}V_{j+1}(z)={}^{\beta}V_{j}(z)$ (61)
    2. (b)
      If there is no conflict because $c_{k}=0$, $a_{k}=\{0,1\}$ and
      1. i.
        if the descend condition is satisfied ($b_{k}=1$) with probability $\rho C(0)(1-B(0))$, then $a_{k}$ descends to $Z_{\mathrm{UP}}$. No UAS enters the $\beta_{Z}$ queue. Following LCFS discipline, the UAS in the $\beta_{Z}$ queue that has arrived last enters service.
        ${}^{\beta}v_{1,k}=0\ ,\ {}^{\beta}V_{1}(z)=1$ (62)
        ${}^{\beta}v_{j+1,k+1}={}^{\beta}(v_{j,k}-1)^{+}$
        $\ \ \ \ \ \ \ \ \ \ \ \ {}^{\beta}V_{j+1}(z)=E[z^{({}^{\beta}v_{j,k}-1)^{+}}]$ (63)
      2. ii.
        If the descend condition is not satisfied ($b_{k}=0$) with probability $\rho C(0)B(0)$, then the arrival $a_{k}$ or the UAS that has arrived last enters service
        ${}^{\beta}v_{1,k}=(a_{k}-1)^{+}=0\ ,\ {}^{\beta}V_{1}(z)=1$ (64)
        ${}^{\beta}v_{j+1,k+1}={}^{\beta}(v_{j,k}+a_{k}-1)^{+}\,$
        $\ \ \ \ \ \ {}^{\beta}V_{j+1}(z)=E[z^{({}^{\beta}v_{j,k}+a_{k}-1)^{+}}]$ (65)
where $E[z^{({}^{\beta}v_{j,k}+a_{k}-1)^{+}}]$ can be obtained from Eqn. 18 and $E[z^{({}^{\beta}v_{j,k}-1)^{+}}]$ can be obtained by substituting $A(z)=1$ in Eqn. 18. The PGF ${}^{\beta}V_{j}(z)$ can be obtained by summing the above conditional PGFs in Eqns. 48-65 using the total probability theorem. By recursively substituting ${}^{\beta}V_{j}(z)$ in ${}^{\beta}V_{j+1}(z)$, we get ${}^{\beta}V_{L}(z)$. In timeslot $k$, a UAS will enter service if there is conflict, otherwise if and only if the system is idle, and more importantly, if there are UAS available to enter service, that is,
@@
    there are UAS in $\gamma_{Z}$ queue (${}^{\gamma}v_{L-1,k}>0$) or,
@@
    there is a nodal arrival ($c_{k}=1$ or $a_{k}=1$, $b_{k}=0$) or,
@@
    there are UAS in $\beta_{Z}$ queue (${}^{\beta}v_{L,k}>0$)
@@
The probability that a UAS will enter the service is given by,
$\sigma_{Z}=A_{\mathrm{I}}(0)\left[{}^{\gamma}V_{L-1}(0)\right]C(0)A(0)\left[{}^{\beta}V_{L}(0)\right]$ (66)
$\pi=(1-A_{\mathrm{I}}(0))+A_{\mathrm{I}}(0){{[}}(1-{}^{\gamma}V_{L-1}(0))$
$\ \ \ \ \ +{}^{\gamma}V_{L-1}(0){\{}{(}(1-C(0))+C(0)(1-A(0))B(0){)}$
$\ \ \ \ \ +C(0){(}(1-B(0))+B(0)A(0){)}(1-{}^{\beta}V_{L}(0)){\}}{{]}}$
$\ \ \ \ \ +\sigma_{Z}(1-A_{\mathrm{O}}(0))$ (67)
$P[w_{1,k}=1]=\Omega_{0}+(1-\Omega_{0})(1-\Theta_{0})\pi\ ,$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ W_{1}(0)=1-P[w_{1,k}=1]$ (68)
@@
where $\sigma_{Z}$ is the descend probability that a non-nodal arrival at node $\widetilde{Z}_{\mathrm{ON}}$ would descend into $Z_{\mathrm{UP}}$ zone and $\pi$ is the probability that a UAS is available to enter service in the given time slot. Note $B(0)=1-\sigma_{\mathrm{I}}$ is the descend probability that the nodal arrival at node $\widetilde{Z}$ will not descend to $Z_{\mathrm{UD}}$ zone.
@@
The Eqn. 68 is the feedback relation that captures congestion responsible for UAS queuing in a Stream $X$ queue. Substituting Eqns. 40-41 in Eqn. 68 and simplifying for $W_{1}(0)$, we get $W_{1}(0)$ as a function of $\Theta_{0}$. Substituting Eqn. 68 in Eqn. 13, we get a polynomial equation in $\Theta_{0}$. Numerically solving the polynomial equation for zeros, we get the congestion probability $\Theta_{0}$. Substituting $\Theta_{0}$ in $\pi$, $\sigma_{Z}$, ${}^{\gamma}V_{j}(z),\ j\in\{1,L-1\}$ and ${}^{\beta}V_{j}(z),\ j\in\{1,L\}$ we get respective probabilities.
As $\theta_{k}$ is an MMRP and $w_{1,k}$ is an MMBP, the respective Markov modulated probabilities $\Theta_{0}^{*}$ and $W_{1}^{*}(0)$ are given by Eqns. 26- 27. The average number of UAS present in the service and queue of Stream $X$ queuing system can be obtained by substituting $\Theta_{0}^{*}$ and $W_{1}^{*}(0)$ in $U(z)$, ${}^{\gamma}V_{L-1,k}$ and ${}^{\beta}V_{L,k}$ and computing $U^{(1)}(1)$, ${}^{\gamma}V^{(1)}_{L-1}(1)$ and ${}^{\beta}V^{(1)}_{L}(1)$, respectively.
@@
The UAS that have queued for $L+1$ slots in the $\beta_{Z}$ queue are lost from the $\overline{Z}$ queuing system and become non-nodal arrivals $a_{\mathrm{I},k}$ for the $\overline{Z}_{\mathrm{ON}}$ queuing system. If the $Z_{\mathrm{UP}}$ is uncongested in time slot $k$, but there are UAS with higher preference present (that is $c_{k}=1$ when there is a conflict or $a_{\mathrm{I},k}=1$ or ${}^{\gamma}v_{L-1,k}>0$ or $a_{k}=1$ when descend condition is unsatisfied or ${}^{\beta}v_{L-1,k}=1$), then in the next timeslot, the UAS corresponding to $q_{k}=1$ is lost from the system. When ZUP turns congested, under steady-state, the fraction of nodal arrivals $a_{k}$ and non-nodal arrivals $a_{\mathrm{I},k}$ queueing in $\gamma_{Z}$ queue that entered the $\beta_{Z}$ queue, but failed to enter service when ZUP was uncongested are lost during the congestion period. The steady-state queue overflow probability $\Phi$ for the $\overline{Z}$ queuing system is given by,
${}^{\beta}v_{L,k}={}^{\beta}v_{L-1,k}+q_{k},\ \mathrm{P}[q_{k}=1]=\left(\cfrac{{}^{\beta}V_{L-1}(0)-{}^{\beta}V_{L}(0)}{{}^{\beta}V_{L-1}(0)}\right)$ (69)
$\Phi=(1-\Theta_{0}^{*}){{{[}}}\ \Omega_{0}(1-C(0))\ +$
${(}C(0)+(1-\Omega_{0})(1-C(0)){)}{{[}}\ (1-A_{\mathrm{I}}(0))\ +$
$A_{\mathrm{I}}(0){\{}\ (1-{}^{\gamma}V_{L-1}(0))+{}^{\gamma}V_{L-1}(0){(}\ (1-A(0))B(0)\ +$
${(}(1-B(0))+B(0)A(0){)}(1-{}^{\beta}V_{L-1}(0))\ {)}{\}}{{]}}{{{]}}}\ \mathrm{P}[q_{k}=1]\ +$
$\Theta_{0}^{*}{{[}}\ (1-A_{I}(0))(1-\gamma V_{L-1}(0))+(1-A(0))B(0)\ {{]}}W_{1}^{*}(0)$ (70)
@@
$A_{\mathrm{I}}(z)=(1-\Phi)+z\Phi$ (71)
IV-C Service departures
Knowing congestion probability $\Theta_{0}^{*}$, the PGF of the probability distribution that UAS would depart the queuing system $\overline{Z}$ after receiving service is given by Eqn. 72. The UAS departures from the system $\overline{Z}$ are $a_{k}$ nodal arrivals for its neighboring upstream system $\overline{Z}_{\mathrm{UP}}$ given by
@@ IV-C Service departures
The algorithmic procedure for estimating the expected spread of UAS traffic in the finite grid $(X,Y),\ X=\{-X_{e},...,0,X_{e}\},\ Y\{1,...,Y_{e}\}$; for a given geometrically distributed request rate $\lambda$ at the source is as follows.
Initialize the bias $\eta$ in traffic about Stream 00 segment.
Initialize the design parameters $L,\ M$.
For zone $Z=(0,1)$ that encloses the source node, initialize
$A(z)=(1-\lambda)+\lambda z,\ A_{\mathrm{O}}(z)=1$
For remaining Level $1$ zones, $Z=(X,1),\ X\neq 0$, initialize
$A(z)=A_{\mathrm{O}}(z)=1$
Start with zone $Z=(X,Y)=(0,1)$
@@
             Evaluate $\Theta_{0}^{*}$ for $\overline{Z}=(X,Y)$ using Eqn. 13;
             Compute average number of UAS in service and queue $U^{(1)}(1)$ and $V_{L}^{(1)}(1)$;
             Compute descend probability $\sigma_{Z}$ for the zone;
             Compute arrivals $A(z)$ for $(X,Y+1)$ zone;
             Compute arrivals $A(z)$ for $(-X,Y+1)$ zone;
             Compute arrivals $A_{\mathrm{I}}(z)$ for $(X+1,Y)$ zone;
             Compute arrivals $A_{\mathrm{I}}(z)$ for $(-X-1,Y)$ zone;
@@
      Set $Y=Y+1$
@@
            Arrival $A_{\mathrm{O}}(z)$ for ($X,Y$) = $A(z)$ of ($X+1,Y$)
             Arrival $A_{\mathrm{O}}(z)$ for ($-X,Y$) = $A(z)$ of ($-X-1,Y$)
@@
Expected spread $\longrightarrow$ Zones where $\Theta_{0}>some\ threshold$
Algorithm 1 Estimating expected spread of UAS traffic
V Theoretical and Simulation Results
Consider the workspace enclosing a source-destination pair at altitude $80\ \mathrm{m}$ in the urban environment. The source is at $(0,0,80)\ \mathrm{m}$, where for every request received, a UAS is deployed to deliver the payload at the paired destination. The time axis is discretized into $\Delta T=2.5\ \mathrm{s}$ time slots. We assume one or no request is received in any given time slot. The average rate $\lambda$ at which the requests are received, analogously the arrival rate $\lambda$ at which the UAS would be deployed at source, varies between $[0.05,1]$ average number of UAS per timeslot. For simulation, a horizontal cross-section of the workspace stretching between [400, $[-400,\ 400]\ \mathrm{m}$ along $x$ axis, [50, $[-50,\ 900]\ \mathrm{m}$ along $y$ axis and intersecting a static obstacle at the given altitude is considered. The UAS is enclosed in a virtual disk of safety radius $r=1\ \mathrm{m}$. The cross-section is tessellated into cells of dimension $5\ \mathrm{m}\times 5\ \mathrm{m}$. The UAS transitions with $1$ cell per time slot nominal velocity. The cell dimensions are large enough that there is no breach in the UAS periphery amongst UAS transitioning in respective cell neighborhoods (shown in Fig. 18). The UAS has a $L=5$ slot look-ahead time window, allowing it to predict where other UAS would be in the next $L\Delta T=12.5\ \mathrm{s}$. Assume that the UTM has prescribed a nominal path between source-destination such that in the cross-section, the set of cells through which the $y$ axis passes is the nominal segment. The cell enclosing the source is the start cell for the nominal segment. The zones are defined with edge length $S=2L+1=11$ cells, and the finite grid (comprising $10$ streams and $10$ levels) is constructed centered on the nominal segment.
Every UAS deployed at the source follows the proposed congestion mitigation strategy. The simulation is run until around $1200$ UAS are deployed at the source. The congestion
probability and expected number of UAS present in the zone are numerically computed by solving Eqn. 13 using MATLAB symbolic math toolbox. The theoretical results are validated with the average values of the $1200$ UAS simulation data.

### v2

@@ 4.2 Stream$\{X\}$ queuing system, $X\neq 0$
        if the descend condition is not satisfied ($b_{k}=0$) with probability $(1-\rho)B(0)(1-\Omega)$, then the UAS that has arrived last enters service.
        ${}^{\gamma}v_{1,k}=(a_{k}+c_{k}-1)^{+}=0\ ,$ ${}^{\gamma}v_{j+1,k+1}={}^{\gamma}(v_{j,k}+a_{k}+c_{k}-1)^{+}$
[…]
The PGF ${}^{\gamma}V_{j}(z)$ can be obtained by summing the above conditional PGFs in Eqns. 40-46 using the total probability theorem. By recursively substituting ${}^{\gamma}V_{j}(z)$ in ${}^{\gamma}V_{j+1}(z)$, we get ${}^{\gamma}V_{L}(z)$. In the timeslot $k$, a UAS will enter service if there is a conflict or if the system is idle and there are UAS available to enter service. In timeslot $k$, a UAS will be available to enter service if,
@@
    there are UAS in $\beta_{Z}$ queue (${}^{\beta}v_{L-1,k}>0$) or,
@@
    there is a nodal arrival ($c_{k}=1$ or $a_{k}=1$ with $b_{k}=0$) or,
@@
    there are UAS in $\gamma_{Z}$ queue (${}^{\gamma}v_{L,k}>0$) or,
@@
The probability $\pi$ that a UAS is available to enter service in the given timeslot is given by,
$\sigma_{Z}=A_{\mathrm{I}}(0)\ {}^{\beta}V_{L-1}(0)\ C(0)\ [1-B(0)(1-A(0))]\ {}^{\gamma}V_{L}(0)\ \ \ \ \ ,\ \ \ \ \pi=(1-\sigma_{Z})+\sigma_{Z}(1-A_{\mathrm{O}}(0))$ (47)
@@
where $\sigma_{Z}$ is the descend probability that a non-nodal arrival at node $\widetilde{Z}_{\mathrm{ON}}$ would descend into $Z_{\mathrm{UP}}$ zone and $1-B(0)=\sigma_{\mathrm{I}}$ is the descend probability that the nodal arrival at node $\widetilde{Z}$ will descend into $Z_{\mathrm{ID}}$ zone. The probability that a UAS will enter the service in a given timeslot is given by,
$P[w_{1,k}=1]=\Omega+(1-\Omega)(1-\Theta_{0})\pi\ ,\ \ \ \ \ \ W_{1}(0)=1-P[w_{1,k}=1]$ (48)
@@
The Eqn. 48 is the feedback relation that captures congestion responsible for UAS queuing in a Stream$\{X\}$ queue. Substituting Eqn. 34 in Eqn. 48 and simplifying for $W_{1}(0)$, we get $W_{1}(0)$ as a function of $\Theta_{0}$. Substituting Eqn. 48 in Eqn. 11, we get a polynomial equation in $\Theta_{0}$. Numerically solving the polynomial equation for zeros, we get the congestion probability $\Theta_{0}$. The existence proof for the $\Theta_{0}$ in $[0,1]$ is similar to the proof of Theorem 1. Substituting $\Theta_{0}$ in $\pi$, $\sigma_{Z}$, ${}^{\beta}V_{j}(z),\ j\in\{1,...,L-1\}$ and ${}^{\gamma}V_{j}(z),\ j\in\{1,...,L\}$, we get the respective probabilities. As $\theta_{k}$ is an MMRP and $w_{1,k}$ is an MMBP, the respective Markov modulated probabilities $\Theta_{0}^{*}$ and $W_{1}^{*}(0)$ are given by Eqns. 20- 21. The average number of UAS present in the service and queue of Stream$\{X\}$ queuing system can be obtained by substituting $\Theta_{0}^{*}$ and $W_{1}^{*}(0)$ in $U(z)$, ${}^{\beta}V_{L-1,k}$ and ${}^{\gamma}V_{L,k}$ and computing $U^{(1)}(1)$, ${}^{\beta}V^{(1)}_{L-1}(1)$ and ${}^{\gamma}V^{(1)}_{L}(1)$, respectively.
@@
The UAS that have queued for $L+1$ slots in the $\gamma_{Z}$ queue are lost from the $\overline{Z}$ queuing system and become non-nodal arrivals $a_{\mathrm{I},k}$ for the $\overline{Z}_{\mathrm{ON}}$ queuing system. If the $Z_{\mathrm{UP}}$ is uncongested in timeslot $k$, but there are UAS with higher preference present that would be entering service, then the UAS corresponding to ${}^{\gamma}q_{k}=1$ is lost from the system. When $Z_{\mathrm{UP}}$ becomes congested, under steady-state, the fraction of non-nodal arrivals $a_{\mathrm{I},k}$ and nodal arrivals $a_{k}$, that end up in the $\gamma_{Z}$ queue, but fail to enter service when $Z_{\mathrm{UP}}$ was uncongested are lost during the congestion period. The steady-state queue overflow probability $\Phi$ for the $\overline{Z}$ queuing system is given by,
${}^{\gamma}v_{L,k}={}^{\gamma}v_{L-1,k}+{}^{\gamma}q_{k}\ \ \ \ ,\ \ \ \ \mathrm{P}[{}^{\gamma}q_{k}=1]=\left(\cfrac{{}^{\gamma}V_{L-1}(0)-{}^{\gamma}V_{L}(0)}{{}^{\gamma}V_{L-1}(0)}\right)$ (49)
${}^{\gamma}\Phi=(1-\Theta_{0}^{*})\left(1-W_{1}^{*}(0)\right)\mathrm{P}[{}^{\gamma}q_{k}=1]+\Theta_{0}^{*}\ (1-A_{\mathrm{I}}(0)+(1-A(0))B(0))\left(1-{}^{\gamma}V_{L}(0)\right)W_{1}^{*}(0)+\xi(1-C(0))$ (50)
@@
$A_{\mathrm{I}}(z)=(1-{}^{\gamma}\Phi)+z\ {}^{\gamma}\Phi$ (51)
4.3 Service departures
Knowing congestion probability $\Theta_{0}^{*}$, the PGF of the probability distribution that UAS would depart the queuing system $\overline{Z}$ after receiving service is given by Eqn. 52. The UAS departures from the system $\overline{Z}$ are $a_{k}$ nodal arrivals for its neighboring upstream system $\overline{Z}_{\mathrm{UP}}$ given by
@@ 4.3 Service departures
The algorithmic procedure for estimating the expected UAS traffic spread in the finite grid $Z=(X,Y)$, where $X=\{-X_{e},...,0,...X_{e}\},\ Y=\{1,...,Y_{e}\}$, for a given UAS arrival rate $\lambda$ at the source is given in Algorithm 1.
For zone $Z=(0,1)$ that encloses the source node, initialize $A(z)=(1-\lambda)+\lambda z,\ A_{\mathrm{O}}(z)=1$ ;
For remaining Level$(1)$ zones, that is $Z=(X,1),\ X\neq 0$, initialize $A(z)=A_{\mathrm{O}}(z)=1$ ;
For zones intersected by exogenous traffic stream, initialize $E_{0}(z)=(1-\lambda_{e})+\lambda_{e}z$ ;
For all zones, initialize the design parameters $L,\ M$ and $\eta$. Start with zone $Z=(X,Y)=(0,1)$ ;
@@
            if for zone $Z=(X,Y)$, the $Z_{\mathrm{UP}}$ is a no-fly zone then
                  Assign $\Theta_{0}^{*}=1$ for queuing system $\overline{Z}=(X,Y)$ ;
             else if for zone $Z=(X,Y)$, the $Z_{\mathrm{IN}}$ or $Z_{\mathrm{ID}}$ is a no-fly zone then
                  Assign $\Theta_{0}^{*}=0$ for queuing system $\overline{Z}=(X,Y)$ ;
             else
                  Evaluate $\Theta_{0}^{*}$ for queuing system $\overline{Z}=(X,Y)$ using Eqn. 11 and Eqn. 20 ;
             end if
            Compute 1) the average number of UAS in service $U^{(1)}(1)$ for zone $Z=(X,Y)$ using Eqn. 8;
                             2) the descend probability $\sigma_{Z}$ for the zone $Z=(X,Y)$ using Eqns. 16, 47;
                             3) the nodal arrivals $A(z)$ for zone $Z=(X,Y+1)$ and $(-X,Y+1)$ using Eqn. 52;
                             4) the non-nodal arrivals $A_{\mathrm{I}}(z)$ for zone $Z=(X+1,Y)$ and $(-X-1,Y)$ using Eqns. 30, 31, 51;
@@
      Expected spread in level $Y$ = Stream$\left[X_{min}\leq X\leq X_{max}\right]$ segments, where for all $X\in\{-X_{e},...,0,...X_{e}\}$,
               $X_{min}=\min\left\{X\ |\ U^{(1)}(1)\geq 1\text{\ for\ }Z=(X,Y)\right\}$ , and $X_{max}=\max\left\{X\ |\ U^{(1)}(1)\geq 1\text{\ for\ }Z=(X,Y)\right\}$;
      Set $Y=Y+1$;
@@
            Arrival $A_{\mathrm{O}}(z)$ for zone ($X,Y$) = $A(z)$ of zone ($X+1,Y$)
             Arrival $A_{\mathrm{O}}(z)$ for zone ($-X,Y$) = $A(z)$ of zone ($-X-1,Y$)
@@
Algorithm 1 Estimating expected UAS traffic stream spread (the set of active parallel stream segments)
5 Simulation Results and Analysis
Consider the workspace enclosing a source-destination pair at altitude $80\ \mathrm{m}$ in the urban environment. The source is at $(0,0,80)\ \mathrm{m}$ where, for every request received, a UAS is deployed to deliver the payload at the paired destination. The time axis is discretized into $\Delta T=2.5\ \mathrm{s}$ timeslots. We assume one or no request is received in any given timeslot. The average rate $\lambda$ at which the requests are received, analogously the arrival rate $\lambda$ at which the UAS would be deployed at source, varies in the range $[0.05,1]$. For simulation, a horizontal cross-section of the workspace measuring [400, $[-400,\ 400]\ \mathrm{m}$ along $x$ axis, and [50, $[-50,\ 900]\ \mathrm{m}$ along $y$ axis and intersecting a static obstacle at the given altitude is considered. The cross-section is tessellated into cells of dimension $5\ \mathrm{m}\times 5\ \mathrm{m}$. The UAS transitions with $1$ cell per timeslot nominal velocity. The UAS is enclosed in a virtual periphery of safety radius $r=1\ \mathrm{m}$. The cell dimensions are large enough that there is no breach in the UAS periphery amongst UAS transitioning in respective cell neighborhoods (shown in Fig. 15). The UAS has a $L=5$ slot look-ahead time window, allowing it to predict where other UAS would be in the next $L\Delta T=12.5\ \mathrm{s}$. Assume that the UTM has prescribed a nominal path between the source-destination pair such that in the cross-section, the set of cells through which the $y$ axis passes is the nominal segment. The cell enclosing the source is the start cell for the nominal segment. The zones are defined with edge length $S=2L+1=11$ cells, and the finite grid (comprising $10$ streams and $10$ levels) is constructed centered on the nominal segment.